7) Use Gauss Law to solve this problem. A 6 nC (nanocoulomb) charge is located at the center of a 2 meter radius uncharged spherical metal shell. Calc. the electric field at a radius of r 1 meter in N/C. (a) 14 (b) 24 (c) 34 (d) 54 8) An electric field is 4 X 10 v/m at an angle of 30 to the x axis. A charge undergoes a displacement vector of 3 m j. (3 meters along the y axis). Calc. the potential difference in volts. (a) 6 X 10% (b) 8 x 10% (c)1x 10 (d) 1.2x 10
Answer #1

(7)

Integral form

E da enc.

Due to spherical symmetry, we have

E(4pi r^2)=rac{Q_{enc.}}{epsilon_0}

Therefore,

9 × 109 × 6 × 10-9 E= = 54 N/C

(8)

V=intvec{E}cdot dvec{S}

E 4 x 10 (cos 307 sin 30), S-3j

Thus,

V=6× 105 V